OpenStudy (anonymous):

Help me out! the sum of twice the first number and three times the second is 10. the sum of one-seventh the first number and 2/3 the second number is -2. What are the numbers?

OpenStudy (anonymous):

let the two numbers be x and y respectively 2x+3y=10 (1/7)x+(2/3)y=-2 then solve these two equations, you can get the ans